OpenStudy (anonymous):

The area of a circle is 28.26 square centimeters. What is its diameter (use 3.14 for ð)?

OpenStudy (anonymous):

area of circle= pi*r^2 28.26=pi*r^2 28.26/3.14=r^2 sqrt9=r 3=r r=3 d=2r 2(3) 6